What is the minimum product of two numbers whose difference is 6?
What are the numbers?

Answer by stanbon(75887)

What is the minimum product of two numbers whose difference is 6?
------
x - y = 6
So x = y+6
-----
product = xy = (y+6)y = y^2+6y
---
You have a quadratic equation:
y^2 + 6y = Product
----
The mimimum occurs when y = -b/(2a) = -6/(2*1)= -3
---
Solve for "x":
x = y+6
x = -3+6 = 3
----
Minimum product = xy = 3*-3 = -9
====================================
Cheers,
Stan H.
====================================